# A three-state coupled Markov switching model for COVID-19 outbreaks across Quebec based on hospital admissions

### v2

@@ 2 A Three-state Coupled Markov Switching Model
Covariates will likely affect the transition probabilities in most cases. For example, dengue outbreaks are unlikely to occur during the winter when mosquito activity is low (Descloux et al., 2012), the introduction of new variants, such as Omicron, has likely played an important role in the development of new COVID-19 waves (Maslo et al., 2022) and disease extinction is less likely in smaller communities (Bartlett, 1957). As for the inclusion of neighboring outbreak indicators in (4)-(5), individuals will mix with those in other areas which can cause outbreaks to spread geographically, a phenomenon known as traveling waves (Grenfell et al., 2001). Therefore, an outbreak in a neighboring area should affect the transition probabilities due to either direct spread or because it indicates spread from a common source, such as a large city.
@@ 5.4 Real-time evaluation and comparison
To evaluate the real-time performance of the models we fit the Full Coupled and Non-coupled models up to 6 weeks before the introduction of Omicron and then up to every week after, that is, we fit the models up to time for , producing 30 sets of posterior samples for each model. We did not include as a covariate since many of the fitted models are fit before, or just after, Omicron and so there is likely not enough information to estimate the effect for many of the fits. Additionally, we found after discarding 27% of the data for the real-time evaluation that MCMC chains would sometimes get stuck in local modes in regions of the parameter space that only used the absence state and one of the count states. This is not an uncommon problem in complex Bayesian mixture modeling and, following the discussion and recommendations in Section 4.2.3 of Frühwirth-Schnatter (2006), we applied mild shrinkage to the transition probabilities by assigning priors to the intercepts in Equations (4)-(5), which stabilized model fitting. Figure 6 compares the posterior probabilities that an outbreak is currently happening (bottom graphs solid lines) between the Full Coupled Model, in orange, and the Non-coupled Model, in purple, for four hospitals. Recall from Section 3.1 that the posterior probability that an outbreak is currently happening is given by and would be used for the purpose of outbreak detection in a real-time scenario. The Full Coupled Model gives an earlier warning of the Omicron outbreak in the first three hospitals, (a)-(c), but not in the final hospital, (d). In (d) the hospitalizations rise very rapidly and so there is a lot of within area information about the Omicron outbreak, making the coupling less useful for providing an early warning. SM Figure 5 (a) shows that, when averaged across all hospitals, early on during the Omicron wave the Full Coupled Model gives a .15-.2 higher posterior probability that an outbreak is currently happening, an earlier warning compared to the Non-coupled Model, while it does not show an increased risk of an outbreak prior to Omicron. The Omicron wave was highly synchronized between the hospitals and, therefore, as can be seen from SM Figure 5 (b), the probability of outbreak emergence would have been higher for the Full Coupled Model during the start of the wave as evidence of outbreaks accumulated in neighboring hospitals. In general, SM Figure 5 shows that the Full Coupled Model tends to enforce the status quo, making outbreak occurrence more (less) likely if there are (are not) outbreaks occurring in neighboring areas.
@@ 6 Discussion
There are also some limitations with our approach. Conditional on the states, we only consider first-order autoregression, see the discussion at the end of SM Section 4. Given weekly data and a disease with a short serial interval, such as COVID-19, this is likely not a major limitation (Bracher and Held, 2022). However, in some cases, especially with daily data, it can be important to account for higher-order temporal dependencies. As another limitation, we assume that the probability of an outbreak emerging does not depend on the amount of time since the last outbreak has ended. After an outbreak ends, the susceptible population should increase over time due to demographic changes, waning immunity and other factors, leading to an increased outbreak risk (Keeling and Rohani, 2007). In a hidden semi-Markov model (Langrock and Zucchini, 2011), the transition probabilities can depend on the amount of time that has been spent in the state, which can be pursued in further work. Finally, given the large number of transitions in our model, the number of covariate effects grows quickly with the number of covariates (4 times). Shrinkage could be applied to the covariate effects in the transition probabilities to eliminate many unimportant effects (Wang et al., 2023).

### v3

@@ 2 A Three-state Coupled Markov Switching Model
Covariates will likely affect the transition probabilities in most cases. For example, dengue outbreaks are unlikely to occur during the winter when mosquito activity is low (Descloux et al., 2012), the introduction of new variants, such as Omicron, has likely played an important role in the development of new COVID-19 waves (Maslo et al., 2022) and disease extinction is less likely in smaller communities (Bartlett, 1957). Note it can often make sense to include the same covariate in both and . For instance, as mentioned, population size should be related to disease extinction, however, it may also affect disease transmission during the endemic and outbreak periods as individuals are more likely to move through high-population areas (Grenfell et al., 2001). As for the inclusion of neighboring outbreak indicators in (4)-(5), individuals will mix with those in other areas which can cause outbreaks to spread geographically, a phenomenon known as traveling waves (Grenfell et al., 2001). Therefore, an outbreak in a neighboring area should affect the transition probabilities due to either direct spread or because it indicates spread from a common source, such as a large city.
@@ 5.4 Real-time evaluation and comparison
To evaluate the real-time performance of the models we fit the Full Coupled, Non-coupled and EE models up to 6 weeks before the introduction of Omicron and then up to every week after, that is, we fit the models up to time for , producing 30 sets of posterior samples for each model. We did not include as a covariate since many of the fitted models are fit before, or just after, Omicron and so there is likely not enough information to estimate the effect for many of the fits. Additionally, we found after discarding 27% of the data for the real-time evaluation that MCMC chains would sometimes get stuck in local modes in regions of the parameter space that only used the absence state and one of the count states. This is not an uncommon problem in complex Bayesian mixture modeling and, following the discussion and recommendations in Section 4.2.3 of Frühwirth-Schnatter (2006), we applied mild shrinkage to the transition probabilities by assigning priors to the intercepts in Equations (4)-(5), which stabilized model fitting. Figure 6 compares the posterior probabilities that an outbreak is currently happening (bottom graphs solid lines) between the Full Coupled Model, in red, and the Non-coupled Model, in blue, for four hospitals. Recall from Section 3.1 that the posterior probability that an outbreak is currently happening is given by and would be used for the purpose of outbreak detection in a real-time scenario. The Full Coupled Model gives an earlier warning of the Omicron outbreak in the first three hospitals, (a)-(c), but not in the final hospital, (d). In (d) the hospitalizations rise very rapidly and so there is a lot of within area information about the Omicron outbreak, making the coupling less useful for providing an early warning. SM Figure 5 (a) shows that, when averaged across all hospitals, early on during the Omicron wave the Full Coupled Model gives a .15-.2 higher posterior probability that an outbreak is currently happening, an earlier warning compared to the Non-coupled Model, while it does not show an increased risk of an outbreak prior to Omicron. The Omicron wave was highly synchronized between the hospitals and, therefore, as can be seen from SM Figure 5 (b), the probability of outbreak emergence would have been higher for the Full Coupled Model during the start of the wave as evidence of outbreaks accumulated in neighboring hospitals. In general, SM Figure 5 shows that the Full Coupled Model tends to enforce the status quo, making outbreak occurrence more (less) likely if there are (are not) outbreaks occurring in neighboring areas.
@@
The top graphs of Figure 6 summarize the one-week ahead predictions, from the previous week, of the hospitalizations from the Full Coupled and Non-coupled models. SM Figure 9 summarizes the one-week ahead predictions from the EE model for the same hospitals. In panels (a)-(c) of Figure 6, there is a much larger discrepancy between the state estimates of the two models than between the predictions of the counts. This indicates that differences in predictive performance may be a poor proxy for differences in the accuracy or, especially, timeliness of the state estimates. Outbreaks often emerge at low counts where there is a lot of overlap between the state distributions, see Figure 5. Therefore, a large difference in the state estimates during the start of an outbreak may not translate to a large difference in predictions. We used the multivariate log score (Bracher and Held, 2022), see SM Section 7.4, to compare the one-week ahead forecasts from the three models. The Full Coupled Model obtained the lowest mean score, indicating the best performance on average across the entire Omicron outbreak, at 3.20. This was followed by the Non-coupled, 3.25, and EE, 3.27, models. Permutation tests (Paul and Held, 2011) showed moderate evidence of a true difference in the overall mean scores, with two-sided p-values of .033, Full Coupled versus Non-coupled, and .052, Full Coupled versus EE. To investigate local differences in forecasting performance, we plot the multivariate log scores across time in SM Figure 10. The Full Coupled Model obtained a much lower score, indicating superior predictive performance, around the emergence of the Omicron wave, which is the most important period to capture for disease surveillance (Buckeridge, 2007). There are some important limitations to using the multivariate log scores to compare the models. We can only evaluate the models over the Omicron wave, where model fitting is stable, which includes just 29 weeks and is missing some important features of the full data set such as long strings of zeroes. In contrast, the WAIC evaluates the models over the entire data set. Finally, an important advantage of the Markov switching models for forecasting, over the EE model, is that they can determine whether a forecasted increase in hospitalizations will likely be due to an outbreak emerging, as opposed to random endemic variation. See the bottom graphs in Figure 6. If an increase in the counts is due to endemic variation it will be much less concerning to health authorities (Unkel et al., 2012).
@@ 6 Discussion
There are also some limitations with our approach. Conditional on the states, we only consider first-order autoregression, see the discussion at the end of SM Section 4. Given weekly data and a disease with a short serial interval, such as COVID-19, this is likely not a major limitation (Bracher and Held, 2022). However, in some cases, especially with daily data, it can be important to account for higher-order temporal dependencies. As another limitation, we assume that the probability of an outbreak emerging does not depend on the amount of time since the last outbreak has ended. After an outbreak ends, the susceptible population should increase over time due to demographic changes, waning immunity, and other factors, leading to an increased outbreak risk (Keeling and Rohani, 2007). In a hidden semi-Markov model (Langrock and Zucchini, 2011), the transition probabilities can depend on the amount of time that has been spent in the state, which can be pursued in further work. Additionally, given the large number of transitions in our model, the number of covariate effects grows quickly with the number of covariates (4 times). Shrinkage could be applied to the covariate effects in the transition probabilities to eliminate many unimportant effects (Wang et al., 2023). Finally, we did not include vaccination data in the model. While data on the percentage of individuals with at least two doses is available in Quebec (INSPQ, 2022), there is very little variation in the covariate. Vaccination coverage increased rapidly and was only greater than zero during the final Omicron outbreak. Also, there are likely interactions between vaccination and new variants (Chenchula et al., 2022). Therefore, we decided there was not enough information in the available data to capture the complicated effects of vaccination with the model. Not including vaccination information could have consequences on the estimated effects of mobility and new variants. For instance, there were fewer mobility restrictions on vaccinated individuals (Cabrera and Olson, 2021) who were also less likely to contract, especially severe, cases of the disease (Johnson, 2022).